Question 1. Find the focal length of a convex mirror whose radius of curvature is 32 cm.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Answer: Radius of curvature, R = 32 cm Radius of curvature = 2 x Focal length (f) R = 2f f = R/2 = 32/2 = 16cm Hence, the focal length of the given convex mirror is 16 cm.